thess is four years younger than bryan the sum of their ages is 60 find their present ages

Answer by srinivas.g(540)

let present ages of Thess and Bryan be x, y respectively
sum of their ages is 60
so x+y= 60
thess is four years younger than bryan
so x+4 =y
put y=x+4 in x+y= 60
x+(x+4)=60
x+x+4=60
2x+4 =60
move 4 to the right
2x=60-4
2x=56
divide with 2 on both sides
2x/2= 56/2
x= 28
y= 28+4 = 32
So ages of thess is 28 years
age of Bryan is 32 years